OpenStudy (jacksonjrb):

Given one or more of its roots, complete the solution set for each equation. Roots Already Given: (0, 5) Need two more.

OpenStudy (jacksonjrb):

\[x^4-\frac{ 13 }{ 2 }x^3+8x^2-\frac{ 5 }{ 2 }x=0\]

OpenStudy (anonymous):

Step one: clear your fractions by multiplying by 2. Step two: multiply the two zeros given. (x)(x-5)= x^2-5x Step three: divide your polynomial by x^2-5x. Step four: your answer in step 3,will be a quadratic equation on which you can use the quadratic equation to find the other two zeros. http://www.youtube.com/watch?v=znsUfxzpiGE